OpenStudy (anonymous):

Find the missing coordinate if A, B, and C are collinear. A = (2, 4), B = (0, 1), C = (-2, y)

OpenStudy (akashdeepdeb):

HINT: If they are collinear then they have the same slope. :)

ganeshie8 (ganeshie8):

or, use segment addition postulate

ganeshie8 (ganeshie8):

if A, B, C are collinear, then AB+BC = AC

OpenStudy (anonymous):

yes.

OpenStudy (akashdeepdeb):

O use area theorem in co-ordinate geometry. :D

OpenStudy (anonymous):

no idea what that is. sorry!! im only a freshman in high school!

ganeshie8 (ganeshie8):

which one u familiar wid ?

OpenStudy (anonymous):

The segment addition.

ganeshie8 (ganeshie8):

that wont be easy here as it gives us quadratic

ganeshie8 (ganeshie8):

quadratic may not be easy to deal wid

OpenStudy (akashdeepdeb):

@hannahbonana Do you know how to find slope? :)

OpenStudy (anonymous):

yes. y=mx+b

OpenStudy (akashdeepdeb):

RISE/RUN? :D

OpenStudy (anonymous):

yes!! y2-y1/x2-x1:)

OpenStudy (akashdeepdeb):

Excellent!! :D Now use that to find the slope from the 2 COMPLETELY GIVEN POINTS!! :D

OpenStudy (anonymous):

ok!! so that would be -3/-2! s0 3/2

OpenStudy (akashdeepdeb):

Yes so m = 3/2 now take any two points B = (0, 1), C = (-2, y) Now even these will have the same slope!! :D Find BC slope and equate it to 3/2!! :D Understood? :)

OpenStudy (anonymous):

uh...you just lost meeee

OpenStudy (akashdeepdeb):

Cna you find the slope of point BC ?

OpenStudy (anonymous):

it would be y-1/-2

OpenStudy (akashdeepdeb):

Or 1-y/2 Now they are collinear right? So the same line has the same slope!! :D 1-y/2 = 3/2 Understood? :)

OpenStudy (akashdeepdeb):



OpenStudy (anonymous):

yes..sort of. so that would be..Y=2!!

OpenStudy (akashdeepdeb):

-2 ? :)

OpenStudy (anonymous):

yes! -y=2 so yes.

OpenStudy (akashdeepdeb):

Understood it? :)

OpenStudy (anonymous):

YES. THANKS!!!!!

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

;D